OpenStudy (anonymous):

Solve : -1/2 x + 3 = −x + 7

OpenStudy (anonymous):

my answer was -2 = -8

OpenStudy (anonymous):

i think i went wrong but where?

OpenStudy (helder_edwin):

add -3 and x to both sides. tell me what u get

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

-1/2 x + 3 = −x + 7 Eliminate the fractions by multiplying the equation by 2 -x + 6 = -2x + 14 add 2x to each side x + 6 = 14 Subtract 6 from each side x = 8